# AURA-CVC: Autonomous Ultrasound-guided Robotic Assistance for Central Venous Catheterization

### v1

@@ 2.5 Needle guidance
The needle insertion parameters, angle ($\theta$) and depth ($d$), are calculated by estimating the kinematic relation between the vein center and remote center-of-motion (RCM, see Fig. 1). The vein center is initially identified in the transverse ultrasound view, but remains at the same pixel location in the sagittal ultrasound view. Let $p^{VC}=[x^{VC},y^{VC},z^{VC}]$ and $p^{RCM}=[x^{RCM},y^{RCM},z^{RCM}]$ be the 3D position of IJ vein center and RCM in robot base frame, respectively. The variables $\theta$ and $d$ can then be estimated using:
$\theta=\tan^{-1}\left(\frac{\boldsymbol{x}^{VC}-\boldsymbol{z}^{RCM}}{|\boldsymbol{x}^{VC}-\boldsymbol{z}^{RCM}|}\right)-\theta_{offset}$ (7)
$d=\|\boldsymbol{p}^{VC}-\boldsymbol{p}^{RCM}\|$ (8)
where, $\theta_{offset}$ is the initial angular offset of the needle (18.21) due to the design constraints of the mechanism. To minimize targeting error, we implemented a closed-loop controller that minimizes the distance between the needle tip and the vein centerline during advancement. As the needle angle couldn’t be adjusted after insertion, the controller was primarily employed to minimize the distance between the needle tip and vein centerline during insertion. The controller used a proportional gain of 0.5 and required manual marking of the vein centerline (one-time) and needle tip (multiple-times) in sagittal ultrasound view after each step. Needle advancement was stopped when the tip is within $1$ $mm$ of the centerline.
@@ 3 Experiments and Results
Table 1: Exp. 1-5 corresponds to scenarios with different orientations (in $deg.$) of phantom with respect to (w.r.t) robot base and Exp. 6-10 corresponds to scenarios with different amounts of deviations (in $mm$) added to the start and end point of the scanning path.
@@ 3.1 Landmark prediction
With 427 test depth images, the mean Euclidean error for 33 landmarks was $13.59$ $mm$$\pm$ $10.49$ $mm$ (95th percentile: $34.26$ $mm$). For the three key CVC landmarks, SN, RLC, and HT, the mean error was $13.88$ $mm$ $11.55$ (95th percentile: $35.35$ $mm$). Table 2 shows detailed results. Although the maximum error reached $34.17$ $mm$, it did not hinder downstream vessel localization, as the triangular scanning region was large enough to encompass the IJ vein in all cases. Nonetheless, reducing this error could improve accuracy in anatomically complex patients.
@@
Figure 7: Comparison of 3D reconstruction of centerline of vein and artery (both marked with magenta color) with their ground truths (blue color for vein and red color for artery) for the 10 experiments
@@ 3.2 Vein-artery centerline reconstruction
We compared the reconstructed centerline with the ground truth center line obtained from CT scans of the phantom. The CT data was processed using the Vascular Modeling Toolkit in 3D Slicer to extract the centerlines discretized at $1$ $mm$ intervals to serve as the ground truth. The reconstructed and ground truth centerlines were aligned using the ICP algorithm (Fig. 7). To assess the accuracy of reconstructed centerlines, we computed their $k-$nearest neighbors distance with the ground truth center-points and results are given in Table 3.
Table 3: Reconstruction accuracy reported as mean, maximum, and standard deviation (S.D.) errors (in $mm$) between reconstructed and ground-truth vein and artery center-lines.
@@ 3.3 Needle targeting accuracy
The targeting accuracy of the system was evaluated by computing the error between the target and final needle tip position. To obtain these measurements, the 3D positions of the targeted point and needle tip were first calculated using eq. (1) from the marked pixel locations in the acquired sagittal view. At the conclusion of each experiment, an expert (co-author VP) evaluated the ultrasound frame and marked the observed needle location and IJ vein (its center point and elliptical boundary). This evaluation was first done in the sagittal ultrasound view after insertion. Then, the ultrasound probe was removed from the robot and manual hand-held scanning was performed to collect the transverse and sagittal ultrasound views from the insertion region. The catheter insertion was also performed to confirm the successful access to the punctured vein. Ultrasound videos were recorded during hand-held scanning and catheter insertion, which were later analyzed by the expert to confirm the safe insertion of the needle and catheter in the IJ Vein for each experiment. Fig. 8 shows the targeted 3D position inside the IJ vein and the final needle tip position after insertion. Table 4 summarizes the corresponding error between the needle tip and vein, both from its center-point and center-line, under the ’Evaluation via Robot’ column.
Fig. 9 shows the annotations by the expert in the selected transverse and sagittal ultrasound frame from the recorded ultrasound video during the hand-held scanning of the region. The transverse ultrasound views show that the needle tip was detected inside the IJ vein boundary. The sagittal views confirm that needle did not puncture through the vein and remained closed to the center line of the IJ vein. Further, the expert confirmed that the catheter was successfully inserted. To evaluate, the error was calculated as the distance between needle tip and center-point in the transverse view of the vein. In the sagittal view, the perpendicular distance between needle tip and center-line of vein was also calculated. These distances were initially recorded in $pixel$ units and later converted to $mm$ using the probe’s scaling factor. Table 4 summarizes these evaluations under the ’Evaluation via expert’ column.
@@ 4 Discussion
We presented experimental phantom results to evaluate vessel centerline reconstruction and needle targeting accuracy. The proposed pipeline and validation study will serve as a foundation for future clinically deployable system. There were also some less-than-ideal performance issues with our experimental goals, which can be explained as follows.
The reconstructed vein exhibited a relatively high mean±$\pm$±S.D. error (Table 3), with two possible explanations. First, the vein has greater compressibility compared to the artery under force-controlled robotic scanning. This compression was not accounted for in the ground truth (i.e., intra-operative CT scan). Second, we attribute this error to the use of convex ultrasound probe. While linear probes are typically preferred for CVC, we only had access to a convex probe for this study. Nonetheless, these results (mean error of vein and artery combined is 2.15 $mm$) indicate that the calibrated transformation from ultrasound image to the 3D space preserves the spatial accuracy of the vein and artery without larger deviation, which is critical for identifying the accurate needle entry point.
The needle targeting accuracy (Fig. 8) shows that the needle tip consistently reached near the IJ vein centerline, with average errors of $0.62$ $mm$ and $0.78$ $mm$ for Exp. 11115555 and 666610101010, respectively (Table 4). These values are close to or under the $1$ mm threshold used in the needle controller (Section 2.5). However, slightly higher average errors from the vein center could be due to bending of the $150$ mm long needle. Still, all errors remained well within the vein’s ground truth radius ($3.24$ to $6.16$ $mm$).
For expert-evaluated manual scans, the mean±$\pm$±SD error from the center-point was $2.49\pm 1.49$ $mm$ and $2.06\pm 0.29$ $mm$ for Exp. 11115555 and 666610101010, respectively. The highest error occurred in Exp. 4 and 5, likely due to unavoidable tilting of the needle mechanism. The mean±$\pm$±SD error from the center-line was $0.63\pm 0.29$ $mm$ and $1.17\pm 0.52$ $mm$, for these two set of experiments, respectively. In Exp. 666610101010, error increased with induced trajectory noise (notably in Exp. 66669999), but decreased again in Exp. 10 due to unpredictable needle bending. Despite such variations, all experiments kept the targeting error within the vein’s radius ($3.24$ to $6.16$ $mm$), confirming effective insertion. Moreover, for contextual reference, Chen et al. [21] reported a positional accuracy of $0.9\pm 0.29$ $mm$ across five needle insertions in a pork phantom.
There are three limitations of the system that can be addressed in future work. First, reducing the needle mechanism’s weight (currently $\sim 4.5$ Kg) and overall size would enhance its positioning flexibility, which would be needed for obese patients. Second, although needle insertion angles were limited by the geometric constraints of the curved rail coupled with the probe, future designs enabling independent motion of the probe and needle could allow for steeper insertion angles, better aligning with clinical preferences. Lastly, manual marking of the needle tip in ultrasound images during insertion limits real-time capability, so future pipeline will explore automated detection.
@@ 5 Conclusion
We have shown that the effective integration of intra-operative planning, anatomical identification, and needle guidance could enable the end-to-end robotic execution of central venous catheterization. To the best of our knowledge, this is the first time a sole depth camera feedback has been used to identify the scanning region, thereby eliminating the need to rely on medical experts or intra-operative CT/MRI scans. Another novelty of this work is the use of force-guided robotic motion to identify the vein and artery based on their compressibility, thereby reducing the reliance on complex ultrasound imaging feedback (multi-class segmentation or Doppler ultrasound). The success of this system can be observed from the successful needle placement with an error below or close to $1$ $mm$. In future, we will use anatomically representative phantoms [26] or fresh cadavers to further validate the system under more realistic physiological conditions.

### v2

@@ 2.5 Needle guidance
The robotic needle mechanism has two degrees of freedom: (i) rotation ($\theta$) to align the needle with the desired insertion angle, and (ii) translation ($d$) to advance the needle to the target depth. These desired angle and depth of needle are computed from the kinematic relationship between the IJ vein center and the RCM of the needle mechanism (see Fig. 1).
Let $p^{VC}=[x^{VC},y^{VC},z^{VC}]$ and $p^{RCM}=[x^{RCM},y^{RCM},z^{RCM}]$ denote the 3D coordinates of the vein center and RCM, respectively, in the robot base frame. The vein center was first identified from the segmented region in the transverse ultrasound plane and then tracked in the sagittal plane. The desired insertion angle $\theta$ is defined as the angle between the line connecting the RCM to the vein center and the vertical $z$-axis of the robot’s end-effector frame, with an offset to account for the mechanism design:
$\theta=\tan^{-1}\left(\frac{x^{VC}-x^{RCM}}{z^{VC}-z^{RCM}}\right)-\theta_{\text{offset}},$ (7)
where $\theta_{\text{offset}}=18.21^{\circ}$ is the initial tilt of the needle guide relative to the robot $z$-axis. The insertion depth is given by the Euclidean distance between the vein center and the RCM:
$d=|p^{VC}-p^{RCM}|.$ (8)
To minimize targeting error, a closed-loop controller was implemented. The controller adjusted the translational motion of the needle to minimize the perpendicular distance between the needle tip and pre-defined vein centerline during advancement. Since the needle orientation was fixed prior to insertion, this control ensured safe puncture and in-vein positioning. The proportional controller used a gain of 0.5. At present, the vein centerline was marked manually once in the sagittal ultrasound view, and the needle tip location was manually identified after each advancement step. The needle was advanced until the tip reached within 1 mm of the centerline.
@@ 3.1 Overview
Table 1: Exp. 1-5 corresponds to scenarios with different orientations (in $deg.$) of phantom with respect to (w.r.t) robot base and Exp. 6-10 corresponds to scenarios with different amounts of deviations (in mmmm) added to the start and end point of the scanning path.
@@ 3.2 Landmark prediction
With 427 test depth images, the mean Euclidean error for 33 landmarks was $13.59$ mmmm $\pm$ $10.49$ mmmm (95th percentile: $34.26$ mmmm). For the three key CVC landmarks, SN, RLC, and HT, the mean error was $13.88$ mmmm ± $11.55$ (95th percentile: $35.35$ mmmm). Table 2 shows detailed results. Although the maximum error reached $34.17$ mmmm, this error did not hinder downstream vessel localization, as the triangular scanning region was large enough to encompass the IJ vein.
@@ 3.3 Vein-artery centerline reconstruction
Fig. 7 shows representative segmentation results of veins and arteries from ultrasound frames. The reconstructed vein and artery centerlines were then obtained by transforming these segmentation outputs in the robot frame using Eq. (1) and compared with the ground truth centerlines extracted from CT scans of the phantom. The CT data was processed using the Vascular Modeling Toolkit in 3D Slicer to extract the centerlines discretized at 11 mmmm intervals to serve as the ground truth. The reconstructed and ground truth centerlines were aligned using the ICP algorithm (Fig. 8). To assess the accuracy of reconstructed centerlines, we computed their $k$-nearest neighbors distance with the ground truth center-point, as reported in Table 3.
Figure 7: Comparison of segmentation network predictions with ground truth on the test set. Predicted regions are shown in blue, ground truth in red, and their overlap in magenta. White boxes highlight zoomed-in views of the overlapping regions.
Figure 8: Comparison of 3D reconstruction of centerline of vein and artery (both marked with yellow color) with their ground truths (blue color for vein and red color for artery) for the 10 experiments
Table 3: Reconstruction accuracy reported as mean, maximum, and standard deviation (S.D.) errors (in mmmm) between reconstructed and ground-truth vein and artery center-lines.
@@ 3.4 Needle targeting accuracy
The targeting accuracy of the system was evaluated by computing the error between the target and final needle tip position. The 3D positions of the target point and needle tip were calculated using Eq. (1) from the marked pixel locations in the sagittal view. After each experiment, an expert (co-author VP) marked the needle location and IJ vein (center point and boundary) in the ultrasound frame. This evaluation was performed first in the sagittal view after insertion. The probe was then removed, and manual handheld scanning collected transverse and sagittal views of the insertion region. Catheter insertion was performed to confirm vein access. Ultrasound videos from hand-held scanning and catheter insertion were later analyzed by the expert to verify safe insertion into the IJ vein.
Fig. 9 shows the targeted 3D vein position and final needle tip location. Table 4 reports the corresponding errors under the ‘Evaluation via Robot’ column. Fig. 10 shows expert annotations in selected transverse and sagittal frames from recorded ultrasound videos. The transverse view confirmed the needle tip remained inside the vein boundary, while the sagittal view re-confirmed that puncture was safe, with the tip remaining close to the centerline. Catheter insertion was also confirmed.
Errors were computed as the distance between needle tip and vein center-point in the transverse view, and as the perpendicular distance to the vein centerline in the sagittal view. These distances were recorded in pixels and later converted to millimeters using the scaling factor described in Section 2.1. Table 4 summarizes these results under the ‘Evaluation via Expert’ column.
@@ 4 Discussion
We presented phantom experiments to evaluate vessel centerline reconstruction and needle targeting accuracy. The proposed pipeline and validation serve as a foundation for a clinically deployable system. Performance was generally acceptable, with three aspects of less-than-ideal performance explained below.
[…]
Second, needle targeting accuracy (Fig. 9) showed tip alignment near the IJ vein centerline, with average errors of $0.62$ mm and $0.78$ mm for Exp. 1155 and 661010 (Table 4). These values fall within the 1 mm threshold used in the controller (Section 2.5). Slightly higher errors arose in Exp. 2, 6 and 9 due to bending of the 150 mm long needle, yet reported values remained within the vein radius (3.243.246.166.16 mm).
[…]
Three limitations should be addressed in future work. First, reducing the footprint of needle mechanism would improve positioning flexibility, which is particularly important for obese patients or those with shorter necks where sagittal access is not possible. Second, independent probe–needle motion could allow steeper insertion angles, better aligning with clinical preferences such as the lateral short-axis insertion emphasized in [rossi2014percutaneous]. Finally, manual marking of the needle tip in ultrasound images currently limits real-time use; a future pipeline will integrate automated needle tip detection and tracking to improve autonomy and reduce operator dependence.
Future work will use anatomically representative phantoms [al2025robotic] or fresh cadavers to validate our system under more realistic physiological conditions. Our immediate next step would be validation of insertion zone selection on human subjects (without insertion). Beyond validation, we also aim to incorporate real-time tissue–needle interaction models and compensation strategies into our framework.
@@ 5 Conclusion
We have shown that the effective integration of intra-operative planning, anatomical identification, and needle guidance could enable the end-to-end robotic execution of central venous catheterization. To the best of our knowledge, this is the first time depth camera feedback has been used to identify the scanning region, thereby eliminating the need to rely on medical experts or intra-operative CT/MRI scans. Another novelty of this work is the use of force-guided robotic motion to identify the vein and artery based on their compressibility, thereby reducing the reliance on complex ultrasound imaging feedback (multi-class segmentation or Doppler ultrasound). The success of this system can be observed from the successful needle placement with an error below or close to 11 mmmm.